# SpeakerLM: End-to-End Versatile Speaker Diarization and Recognition with Multimodal Large Language Models

### v1

@@ Introduction
Impact of Embedding Extractors
In the experiments above, we adopt ERes2NetV2 as the speaker embedding extractor. To investigate the impact of the speaker embeddings, we replace ERes2NetV2 with another speaker embedding model, i.e., CAM++ (Wang et al. 2023), for training and test. It should be noted that ERes2NetV2 performs better than CAM++ in various speaker verification benchmarks (Chen et al. 2025b). Results in Table 4 show that SpeakerLM using embeddings from ERes2NetV2 outperforms the variant with CAM++, indicating that improvements in the speaker embedding model can further enhance the overall performance of SpeakerLM.

### v2

@@ 5.4 Performance with Speaker Registration
5.5 Impact of Embedding Extractors
In the experiments above, we adopt ERes2NetV2 as the speaker embedding extractor. To investigate the impact of the speaker embeddings, we replace ERes2NetV2 with another speaker embedding model, i.e., CAM++ (cam++_2023), for training and test. It should be noted that ERes2NetV2 performs better than CAM++ in various speaker verification benchmarks (3d_speaker_2025). Results in Table 4 show that SpeakerLM using embeddings from ERes2NetV2 outperforms the variant with CAM++, indicating that improvements in the speaker embedding model can further enhance the overall performance of SpeakerLM.